OpenStudy (anonymous):

You buy 3 books that each cost the same amount and a magazine , all for $55.99. You know that the magazine costs $1.99. How much does each book cost? I need solve numerically and algebraically.

OpenStudy (anonymous):

3(X) + 1.99 = 55.99

OpenStudy (anonymous):

Where X = the cost of one book.

OpenStudy (anonymous):

Thanks. What does it mean to solve numerically? Can you show me that?

OpenStudy (anonymous):

Uhmm not a clue, to be honest. Sorry.

OpenStudy (anonymous):

Thanks...no problem. I think it is 55.99-1.99 = 54 54/3=18

OpenStudy (anonymous):

Yeah, that's what I was thinking as well.